????(?? ? ??) = ?(?? ? ??) ? (? ???? ??) ?? =
Answer #1

Proof-

SSxy = \sum [XY-\overline{X}*\overline{Y}]

Opening the summation for the terms inside the bracket-

SSxy = \sum XY-\sum \overline{X}*\overline{Y}

As \overline{X} and \overline{Y} are constants then, \sum \overline{X}*\overline{Y} must be equal to n \overline{X}*\overline{Y}

So, SSxy = \sum XY-n( \overline{X}*\overline{Y})

As \overline{X}=\frac{\sum X}{n}

and \overline{Y}=\frac{\sum Y}{n}

So, SSxy = \sum XY-n( \frac{\sum X}{n}*\frac{\sum Y}{n})

= \sum XY-\frac{\sum X * \sum Y}{n}

Answer

Thanks!